An automobile is accelerating at 5 m/s2 on a straight road on a hill where the radius of curvature of the hill is 200 m. What is the magnitude of the total acceleration when the car’s speed is 30 m/s?


1
Expert's answer
2021-04-26T17:10:03-0400
"a_{total}^2=a^2+\frac{v^4}{r^2}\\a_{total}^2=5^2+\frac{30^4}{200^2}\\a=6.73\frac{m}{s^2}"